OpenStudy (anonymous):

Find dy/dx by implicit differentiation.x^3 + y^3 = 4

OpenStudy (lalaly):

3x^2+3y^2 dy/dx =0 dy/dx =-3x^2 /3y^2 dy/dx= -x^2 / y^2

OpenStudy (anonymous):

yes i knew it thanks!!

OpenStudy (lalaly):

:D